OpenStudy (anonymous):

The rate at which impulses are transmitted along a nerve fiber is directly proportional to the diameter of the fiber. If the rate for a certain fiber is 40 meters per second. When the diameter is 6 micrometers, what is the rate if the diameter is 8 micrometers?

OpenStudy (anonymous):

rate of transmission ∝ diameter of fiber \[40 m/s ∝ 6 µm ⇔ 40 m/s = SomeConstant*6µm⇔SomeConstant = \cfrac{40 m/s}{6 µm}\] \[x m/s ∝ 8 µm ⇔ Diamter m/s = SomeConstant*8µm ⇔SomeConstant = \cfrac{x m/s}{8 µm}\] \[\cfrac{40 m/s}{6 µm} = \cfrac{x m/s}{8 µm}\]